Kohler Corporation reports the following components of stockholders’ equity on December 31, 2016:

Common stock—$20 par value, 100,000 shares authorized,
55,000 shares issued and outstanding

$

1,100,000

Paid-in capital in excess of par value, common stock

80,000

Retained earnings

370,000

Total stockholders' equity

$

1,550,000


In year 2017, the following transactions affected its stockholders’ equity accounts.

Jan.

1

Purchased 4,000 shares of its own stock at $15 cash per share.

Jan.

5

Directors declared a $2 per share cash dividend payable on February 28 to the February 5 stockholders of record.

Feb.

28

Paid the dividend declared on January 5.

July

6

Sold 1,500 of its treasury shares at $19 cash per share.

Aug.

22

Sold 2,500 of its treasury shares at $12 cash per share.

Sept.

5

Directors declared a $2 per share cash dividend payable on October 28 to the September 25 stockholders of record.

Oct.

28

Paid the dividend declared on September 5.

Dec.

31

Closed the $408,000 credit balance (from net income) in the Income Summary account to Retained Earnings.


Required:

1. Prepare journal entries to record each of these transactions for 2017.
2. Prepare a statement of retained earnings for the year ended December 31, 2017.
3. Prepare the stockholders' equity section of the company’s balance sheet as of December 31, 2017.
Answer #1
1
Date General Journal Debit Credit
1-Jan Treasury stock, Common 60,000 =4000*15
Cash 60,000
5-Jan Retained earnings 102,000 =(55000-4000)*2
Common dividend payable 102,000
28-Feb Common dividend payable 102,000
Cash 102,000
6-Jul Cash 28,500 =1500*19
Treasury stock, Common 22,500 =1500*15
Paid-in capital, Treasury stock 6,000
22-Aug Cash 30,000 =2500*12
Paid-in capital, Treasury stock 6,000
Retained earnings 1,500
Treasury stock, Common 37,500 =2500*15
5-Sep Retained earnings 110,000 =55000*2
Common dividend payable 110,000
28-Oct Common dividend payable 110,000
Cash 110,000
31-Dec Income summary 408,000
Retained earnings 408,000
2
KOHLER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2017
Retained earnings, Dec. 31, 2016 370,000
Add: Net income 408,000
778,000
Less: Cash dividends declared (212,000)
Less: Treasury stock reissuances (1,500)
Retained earnings, Dec. 31, 2017 564,500
3
KOHLER CORPORATION
Stockholders' Equity Section of the Balance Sheet
December 31, 2017
Common stock - $20 par value 1,100,000
Paid-in capital in excess of par value, common stock 80,000
Total contributed capital 1,180,000
Retained earnings 564,500
Total stockholders' equity 1,744,500